# Hyperfine and Zeeman Optical Pumping and Transverse Laser Cooling of a Thermal Atomic Beam of Dysprosium Using a Single 421 nm Laser

### v2

@@ Abstract
We demonstrate the effect of Zeeman and hyperfine optical pumping and transverse laser cooling of a dysprosium (Dy) atomic beam on the transition at 421.291 nm. For 163Dy, an electro-optic modulator is used to generate five frequency sidebands required to pump the atoms to the ground state hyperfine level and the light polarization is chosen to pump the atoms to the Zeeman sublevel. The atoms are simultaneously laser-cooled using a standing wave orthogonal to the atomic beam. The resulting polarized and cooled atomic beam will be used in fundamental physics experiments taking advantage of the “accidental” degeneracy of excited states in Dy including the ongoing measurement of parity violation in this system.
@@ I Introduction
Dysprosium (Dy) is an atom of choice in a wide variety of experiments due to its large ground-state magnetic moment (the largest of all atoms), seven stable isotopes with mass numbers ranging from to 164, and the existence of a pair of metastable excited states of the same angular momentum and opposite parity that are nearly degenerate, so that complete degeneracy can be achieved via Zeeman tuning in a magnetic field as low as 1.4 G [1]. These nearly-degenerate levels were used in searches for temporal variation of the fine-structure constant [2, 3] and its dependence on the gravitational potential [4], tests of local Lorentz invariance and Einstein equivalence principle [5], and a search for parity non-conservation [6] (so far focused on 163Dy). It has also been used in searches for apparent oscillation of induced by ultralight scalar dark matter [7, 8].
Dysprosium atoms can be laser cooled [9, 10, 11], trapped [12, 13] and cooled to quantum degeneracy [14, 15], and has become a favored system for the study of quantum magnetism [16].
Much of this work requires knowledge of the spectra of dysprosium, which was studied both theoretically [17, 18] and experimentally [19, 20, 21] in the past few decades. Controlling the internal states of Dy is important for this work, including the ability to efficiently optically pump the atoms to a particular Zeeman sublevel and to a particular hyperfine level in the case of the isotopes with nonzero nuclear spin.
In this work, we use a weakly collimated atomic beam with characteristic transverse velocities of m/s. A collimated narrow-band continuous wave laser only interacts with a small fraction of these atoms. Efficient population transfer of the atomic beam to an excited state requires either broadening the laser linewidth, matching the divergence of the atomic beam and the laser beam [22], or reducing the transverse velocities of the atomic beam by laser cooling. Laser cooling the atomic beam relies on the optical molasses technique, where a standing wave laser beam orthogonal to the atomic beam is used to reduce the transverse Doppler width.
In this paper, we demonstrate such efficient optical pumping and laser cooling using a single laser beam and a custom electro-optic modulator. This is an important step in improving the signal-to-noise ratio in the ongoing search for parity non-conservation in Dy [23].
@@ II Experimental Setup
A schematic of the experimental setup is presented in Fig. 1. A sample of dysprosium is heated up to ∘C to generate the atomic beam. A more detailed description of the source and the UHV setup can be found in Ref. [24]. Two pairs of coils installed inside the optical pumping region along with a pair of coils installed outside the experimental setup were used for the control of magnetic fields inside the optical pumping region. We use the strong 421 nm transition [25] with a lifetime ns and a natural width MHz [26] for optical pumping and laser cooling, and a weaker 598 nm transition [27] ( s; kHz)[21] as a spectroscopic probe.
@@
The 421 nm pump laser is a frequency-doubled master oscillator-power amplifier (MOPA) Toptica DLC-PRO system with an output power of mW. The laser beam is guided as shown in Fig. 1 into the optical pumping region through beam-shaping optics to obtain a rectangular beam with a width (along the direction of the atomic beam) of 20 mm and height of 2 mm leading to a typical interaction time of 65 s. The laser power delivered into the optical pumping/laser cooling region is mW. A part of the fundamental laser frequency output is guided into a Fabry-Perot cavity. The cavity along with the wavemeter reading is used as a reference to lock the laser to Dy transitions.
@@
The probe is a frequency-doubled optical parametric oscillator (OPO) laser (Huebner C-Wave) operating at 598 nm with an output power of 50 mW. After passing through a fiber, and polarization optics, a 1 mm beam diameter laser beam with a power of 1 mW is delivered to the experiment and mW of laser power to a reference photodiode. Both the pump and probe laser beams are circularly polarized to observe hyperfine and Zeeman pumping effects.
@@
The probe laser is scanned across resonance to observe the spectrum of the 598 nm transition of Dy while the pump laser is frequency-locked. Scanning the OPO laser results in periodic oscillatory fluctuations in the laser intensity. The 598 nm fluorescent light is detected using a photomultiplier tube (PMT, Hamamatsu R943-02) oriented orthogonal to the laser and atomic beam, with a 600 nm central wavelength bandpass filter (10 nm FWHM bandwidth, ThorLabs), a colour glass and a homemade spatial filter to reject scattered light from the optical viewports. The PMT signal is fed into a lock-in amplifier and the output of the lock-in amplifier is digitized with an oscilloscope (Tektronix TBS2104) along with the signal from the reference photodiode. The reference signal used for the lock-in comes from an optical chopper (ThorLabs) used for the probe beam. The fluorescence signal is normalized by the signal from the reference photodiode to reduce the effect of laser-power variations.
@@ II.1 Electro-Optic Modulator
A custom broadband electro-optic modulator (QUBIG) is used to generate the frequency sidebands required for the hyperfine pumping of the 163Dy isotope. The pump laser is locked to the 163Dy hyperfine component of the 421 nm transition and five sidebands with frequency detunings of 1190 MHz, 830 MHz, 503 MHz, 235 MHz and 57 MHz are generated [10]. The schematic for optical pumping of these hyperfine and Zeeman states in 163Dy is shown in Fig. 2. The spectrum generated with the EOM is analyzed using a Fabry-Perot spectrum analyzer with a free spectral range of MHz and finesse of .
@@
Figure 3 shows the spectra obtained with the Fabry-Perot spectrum analyzer and the frequency detunings corresponding to all five frequencies being observed. The RF powers used to generate the sidebands are listed in Table 1. A decrease in the laser power in the carrier and sidebands is observed when the EOM is driven at the maximum allowed RF powers of 34 dBm for all modulation frequencies. The 57 MHz modulation is not driven at high RF powers for observation of optimal hyperfine optical pumping due to the strong off-resonant coupling of the pump laser to the hyperfine transition when it is locked to the transition. This off-resonant coupling is observed because of the small frequency shift of 57 MHz () between the two transitions due to hyperfine interactions.
@@ III Experimental results
The frequency of the 421 nm pump laser is locked to a relevant Dy transition for optical pumping and is appropriately detuned to observe optimal simultaneous laser cooling and optical pumping. The fluorescence signal from the atoms along with the reference photodiode signal are recorded as the probe laser is scanned over the isotope and hyperfine components of the 598 nm transition (scan range of 6 GHz; duration of scan 5 s), see Sec.III.3, Fig. 9. Ten different scans are recorded for each spectrum for averaging. The prominent spectral lines are those of 164Dy (natural abundance 28.18 %), 163Dy (24.9 %), 162Dy (25.5 %) and 161Dy (18.9 %). The 163Dy and the 161Dy isotopes both have nuclear spin 5/2 so their spectral lines are split by hyperfine interactions. A bias magnetic field 1.5 G is applied along the direction of propagation of the pump laser beam (chosen as the quantization axis) using the coils inside the optical pumping region.
The reference photodiode signal is used to normalize the fluorescence signal. The position of the 164Dy peak is set as the zero of the -axis and the detuning is calibrated using the isotope shift between the Dy isotopes found in Ref. [27]. The other prominent peaks are identified using the isotope shifts and the hyperfine structure coefficients of the 598 nm transition found in Ref. [27]. Multiple spectra obtained are interpolated to a common detuning and averaged. A nonlinear least squares fit is performed on the normalized and averaged fluorescence spectra. The probe spectrum without the pump laser is fit using a Gaussian profile and the optically pumped spectra are fit with a sum of Gaussian and Voigt profiles. The Gaussian and Voigt profiles are of the form
where and are the Doppler width of the spectrum and width at FWHM due to the laser frequency spectrum, respectively. From Eq. 2, is used to determine the degree of laser cooling. An example of these fits can be found in Fig. 5.
The linewidth of the 163Dy transition is obtained from the fit and the FWHM linewidth is observed to be 32 MHz. The natural linewidth of the transition is 12 kHz [29]. The larger observed width is due to the residual Doppler broadening arising from atomic beam divergence and misalignment of the laser beam, transit-time broadening, and broadening due to the applied magnetic field.
@@ III.1 Polarization Optimization
To optimize the polarization of the 421 nm pump laser, the probe polarization was tuned to induce transitions and the polarization of the pump laser was controlled to induce transitions. The pump laser frequency is tuned to the 164Dy transition. Since the 598 nm probe is a transition, an atomic beam optimally pumped to the ground state in 164Dy would be a dark state for a probe. A vanishing 164Dy spectral line in the probe would indicate optimized pump polarization as seen in Fig 4. This shows optimal Zeeman pumping of the atomic beam but since the atoms are pumped to a dark state of the probe, it is impossible to discern the degree of laser cooling. A probe is used to measure the degree of laser cooling and optical pumping of the atomic beam.
@@ III.2 Zeeman Optical Pumping and Laser Cooling
To observe simultaneous laser cooling and Zeeman optical pumping, the pump laser is first tuned to the 164Dy transition with an appropriate detuning. The increase in the population of the atoms in the , Zeeman sublevel as well as the degree of laser cooling are detected by observing the increase in the amplitude of the spectral line corresponding to the 164Dy probe spectrum. Figure 5 shows the probe spectrum before and after Zeeman pumping and cooling. The 164Dy probe spectrum is amplified by a factor of which is in agreement with the theoretical estimate (Appendix B). The FWHM width of the pumped and cooled spectrum obtained from the fit is MHz. The FWHM of the uncooled probe spectrum is MHz.
The pump laser is then tuned to the 163Dy hyperfine transition and the increase in the amplitude of the spectral line in the probe spectrum is observed. Figure 6 shows the probe spectrum with and without Zeeman pumping and laser cooling the 163Dy isotope. The amplification in the amplitude by a factor of is observed in the probe spectrum. This gain is higher than the expected factor of due to Zeeman pumping because of off-resonant hyperfine pumping from the hyperfine ground state.
The Doppler width of the probe transition at the Doppler cooling limit due to the 421 nm pump is MHz, where is the frequency of the transition and the mass of dysprosium. The Doppler FWHM of the cooled profile of the 163Dy spectrum obtained from the fit to Eq. 2 is MHz and MHz and MHz. The obtained MHz values are of the same order as the expected broadening at the Doppler limit.
@@ III.3 Hyperfine and Zeeman Optical Pumping and Laser Cooling
Figure 7 shows the probe spectrum with and without optically pumping the 163Dy isotope. An amplification of is observed in the hyperfine transition of the probe. This is lower than the theoretical expectation of a factor of 9. The low laser powers in the sidebands Fig. 3 lead to inefficient hyperfine optical pumping.
@@
Accidental hyperfine and Zeeman optical pumping of the 161Dy isotope was also observed with the pump laser tuned to the 163Dy hyperfine transition. This was due to resonances of the frequency sidebands generated by the EOM with the 161Dy hyperfine transitions. Figure 9 shows the amplification in the 161Dy hyperfine transition in the probe spectrum. An amplification factor of was observed in the 161Dy probe spectrum with no laser cooling observed. This is because the closest frequency sideband generated by the EOM is detuned by MHz from the hyperfine component of the 161Dy transition.

### v3

@@ Abstract
We demonstrate the effect of Zeeman and hyperfine optical pumping and transverse laser cooling of a dysprosium (Dy) atomic beam on the transition at 421.291 nm. For 163Dy, a custom built electro-optic modulator is used to generate five frequency sidebands required to pump the atoms to the ground state hyperfine level and the light polarization is chosen to pump the atoms to the Zeeman sublevel. The atoms are simultaneously laser-cooled using a standing wave orthogonal to the atomic beam. The resulting polarized and cooled atomic beam will be used in fundamental physics experiments taking advantage of the “accidental” degeneracy of excited states in Dy including the ongoing measurement of parity violation in this system.
@@ I Introduction
Dysprosium (Dy) is an atom of choice in a wide variety of experiments due to its large ground-state magnetic moment (the largest currently known of all atoms), seven stable isotopes with mass numbers ranging from to 164, and the existence of a pair of metastable excited states of the same angular momentum and opposite parity that are nearly degenerate, so that complete degeneracy can be achieved via Zeeman tuning in a magnetic field as low as 1.4 G [Budker1994_ExpInvDy]. These nearly-degenerate levels were used in searches for temporal variation of the fine-structure constant [Cingoez2007_alpha_var, Leefer2013_alpha_var] and its dependence on the gravitational potential [Ferrel2007_gravitational_dependence_alpha], tests of local Lorentz invariance and Einstein equivalence principle [Hohensee2013_violations_lorentz_symmetry], and a search for parity non-conservation [Nguyen1997_PNC_Dy] (so far focused on 163Dy). It has also been used in searches for apparent oscillation of induced by ultralight scalar dark matter [Tilburg2015_ULDM, Xue2023_UDarkMatter].
Dysprosium atoms can be laser cooled [Muehlbauer2018_optimization_cooling_Dy, Leefer2010_Transverse_cooling, Peterson2019_sawtooth_slowing_Dy], trapped [Youn2010_Dy_MOT, Bloch2023_Dy_tweezer_array] and cooled to quantum degeneracy [Lu2011_BEC_Dy, Lu2012_fermi_gas], and has become a favored system for the study of dipolar quantum gases [Klaus2022_vortices_in_condensate, chomaz2022_dipolar_gases].
Much of this work requires knowledge of the spectra of dysprosium, which was studied both theoretically [Childs1970_ground_state_moments, Dzuba2010_states_Dy] and experimentally [Studer2018_1001nmDy, Budker1991_Spectroscopy_Dy, Schmitt2013_narrow_line_transition_Dy] in the past few decades. Controlling the internal states of Dy is important for this work, including the ability to efficiently optically pump the atoms to a particular Zeeman sublevel [Lecomte2025_production_of_stabilized_dipolar_gas] and to a particular hyperfine level in the case of the isotopes with nonzero nuclear spin.
In this work, we use a weakly collimated atomic beam with characteristic transverse velocities of m/s. A collimated narrow-band continuous wave laser only interacts with a small fraction of these atoms. Efficient population transfer of the atomic beam to an excited state requires either broadening the laser linewidth, matching the divergence of the atomic beam and the laser beam [Nguyen2000_Dy_population_transfer], or reducing the transverse velocities of the atomic beam by laser cooling. Laser cooling the atomic beam relies on the optical molasses technique, where a standing wave laser beam orthogonal to the atomic beam is used to reduce the transverse Doppler width.
In this paper, we demonstrate optical pumping and laser cooling on an atomic beam of Dy, including the 163Dy isotope having well resolved hyperfine structure with energy differences GHz, with a single laser beam and a custom built electro optic modulator (EOM). We also demonstrate “accidental” optical pumping of the 161Dy isotope. The technique adopted in this work can facilitate state preparation in systems with complex hyperfine structure without the need for several repumper laser beams or the use of multiple EOM and acousto-optic modulator (AOM) systems. The results presented in this paper is an important step in improving the signal-to-noise ratio in the ongoing search for parity non-conservation (PNC) in Dy [Leefer2014_towards_PNC].
@@ II Experimental Setup
A schematic of the experimental setup is presented in Fig. 1. A sample of dysprosium is heated up to ∘C to generate the atomic beam. A more detailed description of the source and the UHV setup can be found in Ref. [Weber2013_ac_stark_shifts_Dy]. Two pairs of coils installed inside the optical pumping region along with a pair of coils installed outside the experimental setup were used for the control of magnetic fields inside the optical pumping region. We use the strong 421 nm transition [Leefer2009_Hyperfine_structure_Dy] with a lifetime ns and a natural width MHz [Curry1997_Dy_radiative_lifetimes] for optical pumping and laser cooling, and a weaker 599 nm transition [Childs1977_598nm] ( s; kHz)[NIST_ASD] as a spectroscopic probe. A diagram representing the relevant energy levels for the measurement of optical pumping, laser cooling and PNC in Dy is represented in Fig. 2.
@@
The 421 nm pump laser is a frequency-doubled master oscillator-power amplifier (MOPA) Toptica DLC-PRO system with an output power of mW. The laser beam is guided as shown in Fig. 1 into the optical pumping region through a quarter wave plate and beam-shaping optics to obtain a rectangular beam with a width (along the direction of the atomic beam) of 20 mm and height of 2 mm leading to a typical interaction time of 65 s. The purpose of the rectangular beam shaping is twofold. First, the height is chosen so the laser covers a significant portion of the atomic beam. Second, the length allows for enough pumping cycles to occur ensuring full Zeeman pumping and cooling of the atomic beam (), see Appendix A. The laser power delivered into the optical pumping/laser cooling region is mW. A part of the fundamental laser frequency output is guided into a Fabry-Perot cavity. The laser is locked to the cavity using a side-of-fringe lock and is then software locked to the wavemeter. The wavemeter is then used as a reference to tune the laser to the relevant Dy transition. A thirty minute measurement of the locked pump laser indicates deviations in frequency of MHz.
@@
The probe is a frequency-doubled optical parametric oscillator (OPO) laser (Huebner C-Wave) with a linewidth of MHz operating at 599 nm with an output power of 50 mW. After passing through a fiber, and polarization optics, a 1 mm beam diameter laser beam with a power of 1 mW is delivered to the experiment and mW of laser power to a reference photodiode. Both the pump and probe laser beams are circularly polarized to observe hyperfine and Zeeman pumping effects.
@@
The probe laser is scanned across resonance to observe the spectrum of the 599 nm transition of Dy while the pump laser is frequency-locked. Scanning the OPO laser results in fluctuations in the laser intensity delivered to the atoms and is monitored by the reference photodiode. The scanning probe laser drifts by MHz over a period of thirty minutes which leads to noticeable shifts over scans. The 599 nm fluorescent light is detected using a photomultiplier tube (PMT, Hamamatsu R943-02) oriented orthogonal to the laser and atomic beam, with a 600 nm central wavelength bandpass filter (10 nm FWHM bandwidth, ThorLabs), a colour glass and a homemade spatial filter to reject scattered light from the optical viewports. We measured that 90% of the light collected is transmitted through the coloured glass and bandpass filter setup. The spatial filter results in a 30 dB reduction in the background scattered light with minimal reduction in the atomic fluorescence signal. The solid angle of the collected light is sr. The PMT signal is fed into a lock-in amplifier and the output of the lock-in amplifier is digitized with an oscilloscope (Tektronix TBS2104) along with the signal from the reference photodiode. The reference signal used for the lock-in comes from an optical chopper (ThorLabs) used for the probe beam. The fluorescence signal is normalized by the signal from the reference photodiode to reduce the effect of laser-power variations.
@@ II.1 Electro-Optic Modulator
A custom broadband electro-optic modulator (QUBIG) is used to generate the frequency sidebands required for the hyperfine pumping of the 163Dy isotope following an idea previously proposed in [Leefer2010_Transverse_cooling]. The EOM uses a lithium niobate crystal with an anti-reflection coating for 421 nm, the insertion losses typically range between 0.5 and 0.7 dB. The half-wave voltage of the EOM is V. The RF drivers and mixer were designed and provided along with the EOM by QUBIG. The control unit consists of five RF oscillators tuned to the relevant frequencies, each with their own amplifier. They are then sent to the EOM via a mixer. The pump laser is locked to the 163Dy hyperfine component of the 421 nm transition and five sidebands with frequency detunings of 1190 MHz, 830 MHz, 503 MHz, 235 MHz and 57 MHz are generated. The schematic for optical pumping of these hyperfine and Zeeman states in 163Dy is shown in Fig. 3. The spectrum generated with the EOM is analyzed using a Fabry-Perot spectrum analyzer with a free spectral range of MHz and finesse of .
@@
Fig. 4 shows the spectra obtained with the Fabry-Perot spectrum analyzer and the frequency detunings corresponding to all five frequencies being observed. The RF powers used to generate the sidebands are listed in Table 1. A decrease in the laser power in the carrier and sidebands is observed when the EOM is driven at the maximum allowed RF powers of 34 dBm for all modulation frequencies. The 57 MHz modulation is not driven at high RF powers for observation of optimal hyperfine optical pumping due to the strong off-resonant coupling of the pump laser to the hyperfine transition when it is locked to the transition. This off-resonant coupling is observed because of the small frequency shift of 57 MHz () between the two transitions due to hyperfine interactions.
@@ III Experimental results
The frequency of the 421 nm pump laser is tuned to a relevant Dy transition for optical pumping and is appropriately detuned to observe optimal simultaneous laser cooling and optical pumping. The fluorescence signal from the atoms along with the reference photodiode signal are recorded as the probe laser is scanned over the isotope and hyperfine components of the 599 nm transition (scan range of 6 GHz; duration of scan 5 s); and example of a scan can be found in Sec.III.3, Fig. 10. Ten different scans are recorded for each spectrum for averaging. The prominent spectral lines are those of 164Dy (natural abundance 28.18 %), 163Dy (24.9 %), 162Dy (25.5 %) and 161Dy (18.9 %). The 163Dy and the 161Dy isotopes both have nuclear spin 5/2 so their spectral lines are split by hyperfine interactions. A bias magnetic field 1.5 G is applied along the direction of propagation of the pump laser beam (chosen as the quantization axis) using the coils inside the optical pumping region.
The reference photodiode signal is used to normalize the fluorescence signal. The position of the 164Dy peak is set as the zero of the -axis and the detuning is calibrated using the isotope shift between the Dy isotopes found in Ref. [Childs1977_598nm]. The other prominent peaks are identified using the isotope shifts and the hyperfine structure coefficients of the 599 nm transition found in Ref. [Childs1977_598nm].
The linewidth of the 163Dy transition is obtained from the fit and the FWHM linewidth is observed to be 32 MHz. The natural linewidth of the transition is 12 kHz [NIST_ASD]. The larger observed width is due to the residual Doppler broadening arising from atomic beam divergence which leads to characterized transverse velocities of m/s.
To achieve our goal of efficient simultaneous Zeeman and hyperfine pumping as well as laser cooling of an atomic beam of Dy, we first need to prepare the internal state of the Dy atoms employing the 421 nm laser and the multitone EOM as discussed above and then assess this preparation to optimize it via our experimental parameters. Those parameters include the frequency and polarization state of the 421 nm light and the relative optical power in the sidebands.
@@ III.1 Polarization Optimization
To optimize the polarization of the 421 nm pump laser, the probe polarization was tuned to induce transitions and the polarization of the pump laser is controlled to induce transitions. The pump laser frequency is tuned to the 164Dy transition. Since the 599 nm probe is a transition, an atomic beam optimally pumped to the ground state in 164Dy would be a dark state for a probe. A vanishing 164Dy spectral line in the probe would thus indicate that the pump polarization is successfully set to transfer the ground state population fully into the sublevel. As seen in Fig. 5 this is experimentally achieved by careful alignment of the quarter wave plate. This shows optimal Zeeman pumping of the atomic beam but since the atoms are pumped to a dark state of the probe, it is impossible to discern the degree of laser cooling. Therefore we turn the probe quarter wave plate 90∘ to induce transition and interrogate the population which allows us to measure the degree of laser cooling and optical pumping of the atomic beam.
@@ III.2 Zeeman Optical Pumping and Laser Cooling
To observe simultaneous laser cooling and Zeeman optical pumping, the pump laser is first tuned to the 164Dy transition. An arbitrary detuning of is initially chosen and is experimentally tuned to observe optimal optical pumping. The increase in the population of the atoms in the , Zeeman sublevel as well as the degree of laser cooling are detected by observing the increase in the amplitude of the spectral line corresponding to the 164Dy probe spectrum.
Multiple spectra obtained are interpolated to a common detuning and averaged. A nonlinear least squares fit is performed on the normalized and averaged fluorescence spectra of the well resolved transition. The probe spectrum without the pump laser is fit using a Gaussian profile and the optically pumped spectra are fit with a sum of Gaussian and Voigt profiles. The Gaussian and Voigt profiles are of the form
where and are the Doppler width of the spectrum and width at FWHM due to the laser frequency spectrum, respectively. From Eq. 2, is used to determine the degree of laser cooling. An example of these fits can be found in Fig. 6 which shows the probe spectrum before and after Zeeman pumping and cooling. The 164Dy probe spectrum is amplified by a factor of which is greater than the theoretical estimate (Appendix B) due to laser cooling of the atomic beam. The uncertainty shown here and in the rest of the letter is the uncertainty in the fits. The FWHM width of the pumped and cooled spectrum obtained from the fit is MHz. The FWHM of the uncooled probe spectrum is MHz.
The pump laser is then tuned to the 163Dy hyperfine transition and the increase in the amplitude of the spectral line in the probe spectrum is observed. Fig. 7 shows the probe spectrum with and without Zeeman pumping and laser cooling the 163Dy isotope. The amplification in the amplitude by a factor of is observed in the probe spectrum. Assuming Zeeman pumping of the hyperfine ground state only, we expect an amplification factor of 1.2. The observed gain is higher than the expected one because of off-resonant hyperfine pumping from the hyperfine ground state that brings more population into the hyperfine ground state as well as laser cooling the atomic beam, both of which lead to a larger fluorescence signal. The off-resonant optical pumping from the hyperfine ground occurs due the transition being less than away from the transition.
The expected Doppler width of the probe transition at the Doppler cooling limit due to the 421 nm pump is MHz, where is the frequency of the transition, the mass of dysprosium, and the lifetime of the state. The Doppler FWHM of the cooled profile of the 163Dy spectrum obtained from the fit to Eq. 2 is MHz and MHz and MHz. The obtained MHz values are of the same order as the expected broadening at the Doppler limit. Any residual broadening in the probe spectrum is due to the laser linewidth of 1 MHz.
@@ III.3 Hyperfine and Zeeman Optical Pumping and Laser Cooling
Fig. 8 shows the probe spectrum with and without optically pumping the 163Dy isotope and a comparison of the fits of the Zeeman pumped and hyperfine and Zeeman pumped 163Dy spectrum is shown in Fig. 9. An amplification of is observed in the hyperfine transition of the probe with hyperfine and Zeeman optical pumping. This is in agreement with the theoretical expectation of a factor of 5.4 which indicates that almost all the atoms in the atomic beam have been pumped to the ground state. The optical pumping of atoms in the other hyperfine ground states is also made evident by the lack of spectral lines corresponding to the transitions of the optically pumped hyperfine ground states in Fig. 10.
@@
Accidental hyperfine and Zeeman optical pumping of the 161Dy isotope was also observed with the pump laser tuned to the 163Dy hyperfine transition. This was due to resonances of the frequency sidebands generated by the EOM with the 161Dy hyperfine transitions. Fig. 10 shows the amplification in the 161Dy hyperfine transition in the probe spectrum. An amplification factor of was observed in the 161Dy probe spectrum with no laser cooling observed. This is because the closest frequency sideband generated by the EOM is detuned by MHz from the hyperfine component of the 161Dy transition.